OpenStudy (lexy148):

A jar contains 7 blue cubes, 4 blue spheres, 5 green cubes, and 6 green spheres. if you select an object at random, what is the probability that the object is green or a cube?

jimthompson5910 (jim_thompson5910):

how many cubes are there?

OpenStudy (lexy148):

12 cubes in total

jimthompson5910 (jim_thompson5910):

now how many green objects are there?

OpenStudy (lexy148):

11 green objects

jimthompson5910 (jim_thompson5910):

so we have 12 cubes + 11 green objects = 23 objects that are either green or a cube however, 5 objects are both green and a cube so we subtract this figure out because this is an overlap making us double count: 23 - 5 = 18

jimthompson5910 (jim_thompson5910):

so in reality there are 18 things that are either green, or a cube, or both and we have fixed the double count issue

jimthompson5910 (jim_thompson5910):

now you need to calculate how many objects there are total

jimthompson5910 (jim_thompson5910):

what do you get for that?

OpenStudy (lexy148):

i got 22 for the total

jimthompson5910 (jim_thompson5910):

so the probability is 18/22 = 9/11 I reduced 18/22 to get 9/11. I divided both parts by the GCF 2

jimthompson5910 (jim_thompson5910):

you could do it this way there are 4 objects that are NOT green and NOT a cube (the 4 blue spheres) this is out of 22 objects total, so 22-4 = 18 objects are either green or a cube then do 18/22 = 9/11 to get the same answer

jimthompson5910 (jim_thompson5910):

either way, the final answer is 9/11

OpenStudy (lexy148):

yes i got that now its because i get confused when i read these type of problems

OpenStudy (lexy148):

thank you for the help sorry for wasting your time